Shift report 20011227 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Dec 27 2001 - 03:29:07 EST

  • Next message: Apache: "Shift report 20011227 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Dec 27 03:29:07 2001, Shift: 16:00-24:00
    Supervisor: fv, with: 
    Log book#11 pages 96 - 97
    
    Shift Summary:
    ------------------------------------------------------------
    Beam off, when I came in (store aborted at ~15.30)
    
    16:35 Snake quenchlink in sector 3. Estimated time for injection 17:30
    
    Beam studies scheduled for tomorrow Dec 27 (8-14) mainly at injection energy.
    Shift will be on standby until noon.
    
    Plan for shift:
    
    a) Add rhic clock (blue of yellow) to Inel trigger
    b) check effect of adding in theneartest two (L1 and R1) to trigger
       (should improve eff by ~ *2, and according to pp2pp only add a few 
       percent of wrong
    
    
    16:55 Did one more attempt to turn on/off the 1458B. It still gives the
          Bad 24V. Thermal overload ? Turned off until access can be made.
    
    17:05 HV check performed.
    
    18:20 Beam cogged and stored. (99 GeV)
    
    18:40 starting physics run.
          Run 6157  - scaledown 100, 100, 1, 1.
    
    18:50 The trigger six rates does not look right. Somehow there are more 3's than 6'
          Discard run 6157- bad timing!!
    
    19:15 start run 6158.  (8 deg 1/8 B no BFS, 90 deg 350A - B pol.)
          Blue life time is not good.
          Coin rate i.e. Inel decya's rapidly ~500 at 19.50
          The spec eff is smaller for this store ~ 777/1800 (40%) for FS
          Trigger 3/trigger6 ~ 1/3.
    
    20:15 Stop run 6158 - MCR called for start of pol. measurements.
          Will use time to look if L1 and R1 can be added into trigger (for higher eff
          trigger)
    
    20:40 Polarization measurements completed.
         	MCR called plan to hold store until ~ 23.00 
    	where STAR+Phenix has requested access.	
    20:45 Run 6159 started
          down scale 1: 100 5 :20 3 and 6 : 1
    
    20:47 Lpd print server was stopped. - restarted (as su)
    
    22:35 MCR announced another pol measurement.
          Stop data run.6159
    
    22:40 Start run 6160
          Checkout of data quality during polarization measurements.
          The run stopped at 23:05. Should not be used for physics analysis
          Adjusting timing during this run.
          
    
    23:05 Finished sdjusted the timing for trigger 6 start.
          It was quite late ~ 24 nsec.
    
    23:06 Start run 6161
          same trigger mix as before.
    
    23:10 Announchment made for starting access at midnight (1 (2?) hours)
          Beam will go onto 23:55
    
    Run summary:
         Run 6158  tr1-28.6K tr3 - 2914K tr 5 47.8 K tr 6 10271 (1.0 h)
         Run 6159  tr1 66.8K tr 3  3717  tr 5 212K   tr 6: 13613 (1:50h)
         Run 6161  still going at midnight
    
         8 deg  1/8B  ~50k trigger 6 (get to 300K)
         90 deg 2B    ~10k trigger 3 (get to 300K)
    
    Beam Summary:
         18:40-19:15   Bad run (trigger problem)
         19:15-20:15   Data run 
         20:45 - 22:40 Data run
         23:05 - 24:00     Data run
    
    Accomplishments
        - The rhic clock was added to the Inel trigger.
        - trigger 6 timing adjusted.
        - The gitter between different L1.R1 makes it difficult to include until
          we have TDC information for all channels.
    
    Open Work to do:
        schedule access for HV repair (RAS+..- request 2-3 hours)
        string the Inel counter tdc delay cables (ready on spool in FEH).
    
    Beam Hour this shift (16-24)
         4.5 h 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Dec 27 2001 - 03:29:48 EST